OpenStudy (anonymous):

What is the equation of the line that passes through (4, 3) and (-4, -1)? 2x + 1 y = 2x + 7 y = + 1 y = + 7

OpenStudy (anonymous):

@peachpi ???

OpenStudy (anonymous):

do you know the formula for slope?

OpenStudy (anonymous):

is it y2-y1 over x2-x1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so do I just do that to find out the answer?

OpenStudy (anonymous):

it looks like you actually have to write the equation. After you find the slope put it and one of the points into point-slope form. Then you solve for y

OpenStudy (anonymous):

y - y1 = m(x - x1) is point slope

OpenStudy (anonymous):

Is the answer y=x/2+1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you for your time kind sir

OpenStudy (anonymous):

you're welcome, although I'm not a sir. Just like Ernie :)